OpenStudy (anonymous):

Solve the following for x l (x+h) ^2 - x^2 l > or equal to 3h^2, h>0

OpenStudy (anonymous):

@wio could you help me with this problem?

OpenStudy (anonymous):

\[ |(x+h)^2-x^2|\geq 3h^2 \]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First simplify inside: \[ (x+h)^2-x^2 = x^2+2hx+h^2-x^2 = 2hx+h^2 \]

OpenStudy (anonymous):

Next, we know \(h>0\implies 3h^2>0\) so both are positive, meaning if we square them then the inequality will remain true.

OpenStudy (anonymous):

\[ (2hx+h^2)^2 \geq 9h^4 \]

OpenStudy (anonymous):

now do I just simplify that and solve for x?

OpenStudy (anonymous):

Yeah, you wanna solve for the \(=0\) case. You can use quadratic formula.

OpenStudy (anonymous):

When you know the points where it crosses the \(0\) mark, you know when the equality changes between true and false.

OpenStudy (anonymous):

You'll end up with something along the lines of \[ (x-r_1(h)) (x-r_2(h)) \geq 0 \]Where \(r_1\) and \(r_2\) are roots, and they happen to be functions of \(h\).

OpenStudy (anonymous):

oh okay got it, so do I do \[\left| x-1 \right|+\left| x-2 \right|<3 hint: consider \cases for x \le 1, 1<x \le 2, x>2\]

OpenStudy (anonymous):

??? No idea where you got that.

OpenStudy (anonymous):

sorry I meant to say, would I do that problem the same way I did the other one?

OpenStudy (anonymous):

that was a miswording from my side

OpenStudy (anonymous):

No, the other problem is completely different.

OpenStudy (anonymous):

ohh could you please explain how it's different? Just to know for future purposes

OpenStudy (anonymous):

It's not a quadratic equation.

OpenStudy (anonymous):

ohh how would approach this problem then?

OpenStudy (anonymous):

but basically, you have to consider when \(|x-2|>0\) and \(|x-3|>0\)

OpenStudy (anonymous):

that is why they said to consider \(x>2\) and \(x>3\) and also \(x\lt 2\)

OpenStudy (anonymous):

whoops... should be \(1\) and \(2\)... not \(2\) and \(3\).